solve.
the sum of two numbers is 12, and twice one number minus the other is -4. what are the numbers?
heres my equations.
x+y = 12 , x * 2-y = -4
show all work please

Answer by Alan3354(69443)

the sum of two numbers is 12, and twice one number minus the other is -4. what are the numbers?
heres my equations.
x+y = 12 , x * 2-y = -4
--------------
You have the right equations, unless it's 2*(x-y) = -4. Some ambiguity there.
x + y = 12
2x - y = -4
------------- Add
3x = 8
x = 8/3
y = 28/3